### v1

# SELF-EXPLAIN: Teaching Large Language Models to Reason Complex Questions by Themselves

Abstract
Large language models (LLMs) can generate intermediate reasoning steps. To elicit the reliable reasoning, the common practice is to employ few-shot chain-of-thought prompting, where several in-context demonstrations for reasoning are prepended to the question. However, such chain-of-thought examples are expensive to craft, especially for professional domains, and can have high variance depending on human annotators. Therefore, this work investigates whether LLMs can teach themselves to reason without human-crafted demonstrations. We propose SELF-EXPLAIN to generate CoT examples by LLMs inspired by “encoding specificity” in human memory retrieval. We find using self-explanations makes LLMs more confident, more calibrated and less biased when answering complex questions. Moreover, we find prompting with self-explanations can even significantly outperform using human-crafted CoTs on several complex question-answering datasets.
1 Introduction
Recently, with the appearance of large langauge models (LLMs), some sparks of human-like intelligence are shown. LLMs can learn the underlying patterns from several in-context exemplars, which is known as in-context learning (ICL) (1). In ICL, no weight updates are implemented and LLMs construct answers based on their encoded knowledge following the given in-context demonstrations. On the other hand, Chain-of-Thought (CoT) prompting (31) shows that LLMs can generate rationale while answering the questions. Moreover, not only can LLMs learn to reason from in-context CoT examples, but also generate explanations in zero-shot prompting (13). However, despite the evidence of zero-shot reasoning ability, LLMs generally perform badly without any in-context demonstrations, especially when the tasks are complex. High-quality human-crafted CoT demonstrations, which are usually considered as the gold standard (17), are needed to guide LLMs to reason correctly through ICL.
However, designing and collecting CoT demonstrations for prompting LLMs can be expensive and difficult. On the one hand, for some professional domains such as medicine, physicians or medical students may be needed to produce fine-grained explanations of the answers (20; 32). On the other hand, collecting CoT examples (usually in the form of long sentences) through crowdsourcing can have great uncertainty and noise (29; 7). The style or the form of CoTs may mainly depend on human annotators’ own experience. However, Coda-Forno et al. (3) show that different kinds of expressions of context can actually lead to distinct behaviors of LLMs. Additionally, it is difficult to measure whether those human-crafted CoT examples inform the correct distribution and feature space of the task-specific corpus for effective prompting (18).
Therefore, we are motivated to investigate whether LLMs can teach themselves to reason complex tasks without relying on human-designed CoTs. Our contributions are mainly in three folds: (1) We find LLMs themselves can produce explanations of complex questions given the answer, which we refer to SELF-EXPLAIN. (2) Prompting with self-explanation  [^1] We use CoT and explanation interchangeably to refer to sentences showing reasoning in the prompt. through ICL also makes LLMs more calibrated and less biased. For example, our experiments show self-explanation makes LLM 1.2% less biased in multiple-choice questions. (3) Prompting with self-explanation can outperform prompting with human-crafted CoT on challenging reasoning tasks even for clinical QA which requires professional medical knowledge. Our findings may challenge the common wisdom that human-crafted CoTs are the gold standard for prompting LLMs.
2 Related Work
In-Context Learning.
In-Contex Learning (ICL) is the ability of language models to induce answers from given demonstrations without weights updating in supervised tuning. In-context exemplars are the key to ICL which have dominating influence on the generation. Quite a few works have been proposed to optimize the selection of exemplars (16; 24; 4). On the other hand, in the cases of no access to task labels, Lyu et al. (18) proposed zero-shot ICL that employs pseudo exemplars with random labels for classification tasks. Such zero-shot method can be supported by the finding of Min et al. (19) on classification tasks that the label space carries more weight than correctness of labels in exemplars for ICL.
Chain-of-Thought Prompting.
Prompting with reasoning triggers LLMs to generate intermediate steps of thinking, known as Chain-of-Thought (CoT) (31). Some works (22; 12; 21; 35) on multi-step reasoning prompt LLM to decompose a complex question into sub-tasks iteratively, showcasing the LLMs’ capability of compositional reasoning under some human guidance. Additionally, Kojima et al. (13) propose zero-shot prompting to elicit LLMs’ chain-of-thought reasoning without any human-crafted exemplars. This prompting method is leveraged to generate CoT exemplars for ICL (28; 33; 2). Wan et al. (28); Zhang et al. (33) mainly focus on selecting CoTs generated through Kojima et al. (13). Chen et al. (2) further incorporate pseudo task generation apart from pseudo label/ CoTs. Different from those works, we explore a new zero-shot prompting fashion inspired by encoding specificity in human’s memory retrieval. Additionally, we further rigorously investigate the difference between human-crafted CoTs and self-generated CoTs and the according influence on model confidence.
3 Revisiting ICL
We first detail some annotations and give a formal setup of ICL. We denote the model parameters as $\theta$, CoT as $\pi$ and assume a labeled dataset $\mathcal{D}$ with distribution $p^{\ast}$. Given a test query $\mathbf{x}$, the model will predict $\tilde{\mathbf{y}}$ by conditioning on the query and in-context exemplars. We can then have,
$\tilde{\mathbf{y}}=\text{argmax}_{y}P(y|\mathbf{e},\mathbf{x},\theta),$ (1)
where $\mathbf{e}$, is the sequence of all $K$ in-context exemplars i.e., $\mathbf{e}=e_{1},...,e_{K}$) and $e_{i}=(x_{i},\pi_{i},y_{i})$ where $(x_{i},y_{i})$ is sampled from $p^{\ast}$.
3.1 Encoding Specificity Hypothesis
A key question for CoT prompting is how to compose in-context CoTs for some specific dataset $\mathcal{D}_{(x,y)}$? CoTs can be rephrased differently while delivering the same logic. To understand this question, we take a memory view of ICL by conceptualizing LLMs as memory networks (6; 11; 34; 10; 23; 14). The feed-forwarding through hidden layers of LLM is to retrieve and generalize learnt information in memory to construct the output $\mathbf{y}$ to complete the query $\mathbf{x}$ under the guidance of context $\mathbf{C}$ (i.e., in-context exemplars). The pretraining stage can be viewed as encoding information into the weights, i.e., memories of LLM.
From a memory view, CoTs should be composed in a way to match the context in the training corpus seen during encoding similar information to the test data. Such requirement on the match of context between testing and training is known as encoding specificity (27) in humans’ memory retrieval. To see this hypothesis, a simple thought experiment can be considered: supposing that a specific datapoint $(x,y)$ has been seen during language modeling in pretraining and $C$ is the corresponding context prepending $(x,y)$, (i.e., a consecutive string $(C,x,y)$ is seen by LLM during training), at test time, prompting the LLM with $(C,x)$ will trigger the LLM to generate $y$ with great probability. All in all, the encoding specificity hypothesis implies in-context exemplars should match the distribution of context seen during pretraining, i.e., the training corpus encoded by LLMs containing similar information to $(x,y)$. In that way, LLMs are expected to be motivated to search from related encoded knowledge to form the output generation.
Figure 1: The overall framework of SELF-EXPLAIN. We first generate self-explanations on some sampled training data. At test time, those self-explanations will be used as in-context exemplars for prompting.
4 SELF-EXPLAIN
Motivated by the encoding specificity, we would like our CoTs to match the training corpus similar to the test data distribution. In order to achieve this, we propose SELF-EXPLAIN where we directly prompt the LLM to elicit the explanation. The overall pipeline of our approach is shown in Fig. 1. The LLM will leverage the encoded knowledge it deems related to the provided question so as to finish the according generation. Those self-explanations are then used as in-context CoT exemplars for ICL. Formally, we consider a setting where we have some data $(x,y)$ sampled from the distribution $p^{\ast}$ and $p^{\text{test}}\approx p^{\ast}$. We show $y$ does not have to be ground-truth in Sec. 5.2. We assume no human-crafted CoTs are available and define the best CoT as $\pi^{\ast}:=\text{argmax}_{\pi}P(y|x,\pi,\theta)$. Self-explanation is represented as,
$\pi^{\text{self}}=\text{argmax}_{\pi}P(\pi|x,y,\gamma,\theta),$ (2)
where $\gamma$ is an instruction. We hope LLMs to generate reasoning path based on given $(x,y)$ by recalling from its according encoded knowledge so as to satisfy encoding specificity. Moreover, we find $P(y|x,\pi^{\text{self}},\theta)>>P(y|x,\pi^{\text{human}},\theta)$ and $P(y|x,\pi^{\text{self}},\theta)\approx 1$. Further discussion with supporting experimental results is in Sec. 6. We may arguably state that $\pi^{\text{self}}$ is a very close estimation to $\pi^{\ast}$ and have the following approximation,
$\text{argmax}_{\pi}P(\pi|x,y,\gamma,\theta)\approx\text{argmax}_{\pi}P(y|x,\pi,\theta).$ (3)
This shows the self-explanation to a specific $x$, is actually more effective and acceptable for LLM to generate the according $y$, which may evidence the encoding specificity hypothesis. Because self-explanation may be generalized from some related training data in a similar distribution to $x$.
Generalization through generation diversity.
Generated self-explanation will then be used as in-context demonstrations together with corresponding question-answer pairs. However, the underlying logic of $\pi^{\text{self}}$ may be very specific to its corresponding $(x,y)$ and thus lacks generalizability to other different cases. Then, the output explanation $\hat{\pi}_{\text{te}}$ at test time may fail to apply to the input cases, leading to wrong answers. To mitigate this issue, we design $\gamma$ so as to prompt the model to generate solutions employing distinct logics. Formally, we have,
$(\pi^{\text{self}}_{1},...,\pi^{\text{self}}_{n})=\text{argmax}_{\pi}P(\pi|x,y,\gamma,\theta),$ (4)
where $n\in(1,N)$ and NN is the number of different explanations to generate. Then $\pi^{\text{self}}_{i}$ for some $(x_{i},y_{i})$ as an in-context exemplar will be randomly sampled from the according $\{\pi^{\text{self}}_{n}|n\in(1,N)\}$ of $(x_{i},y_{i})$ for unseen test cases.
5 Experiments
5.1 Experimental Setup
We are focused on knowledge-intensive question-answering tasks that require logical thinking on information and associating encoded knowledge but mere facts retrieval. Such knowledge-intensive QA is common and important for the applications of LMs (9; 26). We use challenging MedMCQA (20) and MedQA (8) with standard splits that involve multiple-choice questions to diagnose clinical cases, which are used for physician qualification exams. We also evaluate SELF-EXPLAIN on StrategyQA (5) for reasoning in general domains. In terms of implementations, we use GPT-3.5-turbo as our base LLM. Except for Auto-CoT (33), we randomly sample five exemplars from training data and keep exemplars the same for different baselines using ICL. For MedQA which does not provide human-crafted CoTs, in all cases, we use all the five exemplars provided and annotated in Singhal et al. (25).
[…]
No CoT 51.7 55.1 46.8
Zero-shot CoT (13) 51.1 60.6 45.6
Auto-CoT (33) 52.5 61.9 52.7
Human CoT 53.1 61.7 56.1
Self-Exp 56.6 63.7 59.7
Table 1: Test accuracy of different prompting methods on three datasets.
5.2 Results
As is shown in Table 1, using CoT prompting gives better performance than using input-output pairs solely. This demonstrates the importance of CoT in complex reasoning tasks to LLMs. More importantly, prompting with our proposed self-explanation can impressively outperform using CoTs crafted by human professionals by 2% in challenging Medical datasets. Considering the difficulty and expense of crafting CoTs by humans, employing SELF-EXPLAIN can thus be very useful in some domains. Our approach also outperforms Auto-CoT (33) and vanilla zero-shot CoT (13). Additionally, our results may challenge the common belief that human-crafted CoTs are gold standards which are expected to give better performance than using machine-generated CoTs.
MedMCQA MedQA StrategyQA
Right 56.6 63.7 59.7
Wrong 56.0 63.5 59.1
Table 2: Test accuracy of prompting with self-explanations that are generated provided by right answers and wrong answers.
Does the correctness of self-explanation matter?
A natural question raised in SELF-EXPLAIN is what if the self-generated CoTs are wrong since the generation process is not supervised by humans. We look into this question by guiding LLMs to generate misleading explanation. More specifically, we provide the LLM with wrong answers to generate corresponding self-explanations. Those self-explanation for wrong answers are then used for prompting as in-context exemplars. The results are shown in Table 2. We find that the performance of prompting with self-explanation is insensitive to its correctness. We speculate that what carries more weight is the underlying distribution of self-explanations. This may evidence encoding specificity hypothesis that the matched distribution of exemplars at test with that of the context seen during encoding relevant information at pre-training is more important for effective prompting through ICL. This result suggests that a labeled dataset is not necessary to elicit CoTs in SELF-EXPLAIN.
Figure 2: Similarity between human-crafted CoTs and self-explanations in terms of ROUGE-L score, terminology covered (terms are extracted by SciSpacy) and length on MedCQA.
How similar are self-explanations and human-crafted CoTs?
The common standard to measure the quality of machine-generated samples is how similar they are to human-crafted ones (17; 30). The more similar, the better the quality is. However, despite the superior performance of self-explanation in prompting, we find that they are very different from human-crafted CoTs in terms of ROUGE-L score, terminology used and length. We use MedCQA as testbed which provides high-quality human-crafted explanations. For terminology comparison, we extract terms in both kinds of CoTs through scispaCy  [^2] https://allenai.github.io/scispacy/ and calculate F1 score between the two terms lists. Results are shown in Fig. 2. We find in terms of content (measured by ROUGE-L and term coverage), self-explanation differs from human-crafted CoTs greatly given the average similarity is around 15%. The length distribution of self-explanation is more centric, while the human-crafted CoTs have more varied lengths. We attach some example self-explanations in Appendix D.
6 Influence on Model Confidence
Figure 3: Comparison of model confidence in different cases by using human-crafted CoT and self-explanation. Prompting with self-explanations makes the model more calibrated and more confident as well. (a) Comparison between confidence of true answers and false ones. (b) Further comparison of confidence in more specific cases. “S” denotes “Self_exp”, “H” denotes “Human_cot”, “T” denotes “True” and “F” denotes “False”, e.g., “S_T & H_F” means self explanation gives true answer, while human-crafted CoT leads to false answer.
Figure 4: Average confidence of model on each option for a balanced dataset where the number of questions that each option is the correct answer is equal. The correct options in exemplars are fixed to the same choice in each sub-figure for more rigorous comparison. The selection bias is calculated as $\frac{1}{4}{\sum_{i=1}^{4}|q_{i}-0.25|}$ where $q_{i}$ is model’s confidence for each option.
Models are more confident when prompted by self-explanation.
For exemplars selected for generating self-explanation, we use Text-Davinci-003 to compute the average $P(y|x,\pi^{\text{self}},\theta)$ which reaches 99.96%, which may support our estimation of Eq. 3 in Sec. 4. In comparison, for human-crafted explanation of the same exemplars, average $P(y|x,\pi^{\text{human}},\theta)$ is lower, reaching 89.05%. This implies that given $(x,y)$, self-explanation may be the most related context to elicit $y$. For inference with self-explanation as demonstrations in ICL, log probabilities are computed for correct and wrong model outputs prompted with self-explanation and human-crafted one. Results are shown in Fig. 3. We can observe that models’ output log probabilities with self-explanation are much higher than with human-crafted explanation, indicating greater model’s confidence in its output. This suggests self-explanation can be more acceptable and effective for LLMs to elicit reasoning.
We next consider two metrics measuring the influence of prompting with self-explanation and human-crafted CoT on LLM. (1) Calibration: the model’s output confidence should indicate the correctness of answers (e.g., wrong answers have lower confidence). Well-calibrated output confidence can assist human users to determine whether to trust model’s outputs or to look for alternatives in high-stakes applications. (2) Selection bias: the model should give equal chance of each option for a balanced multi-choice QA dataset. The experiment implementations are detailed in Appendix C. We find that (1) Self-explanation improves calibration. Fig. 3(a) showcases the model is generally more calibrated when using self-explanation. Further results in Fig. 3(b) show that output confidence by using human CoT is not indicative especially when the question is debatable (i.e., using self-explanation and human CoT generate different answers). Such results are consistent among different LLMs (see Fig. 5 in Appendix). (2) Self-explanation reduces bias. We control the correct option of in-context exemplars to be the same (e.g., all the correct options are A) for rigorous comparison. Results in Fig. 4 show that in all cases, using human-crafted CoTs leads to higher selection bias than using self-explanations.
7 Conclusion
In this work, we delve into whether LLMs can teach themselves to reason complex questions. Chain-of-Thought (CoT) can be costly to craft by humans and may have great uncertainty and noise. Therefore, we propose SELF-EXPLAIN inspired by encoding specificity hypothesis in memory retrieval to prompt LLMs to elicit reasonings given a question-answer pair. We find that when doing few-shot prompting with self-explanations as in-context CoT demonstrations, the LLM can surprisingly reach better performance than prompting with human-crafted CoTs on complex questions that require reasoning. We also find that LLM is more calibrated and less biased when prompting with self-explanations. However, comparison on four metrics shows that self-explanation is very different from human-crafted CoTs. Our work may thus challenge the common wisdom that human-crafted CoTs are the gold standard for prompting and possibly shed new light on understanding the effects of exemplars on ICL.
Appendix A Effects of Cues in SELF-EXPLAIN
In this section, we examine the performance of ICL with self-explanations prompted by different cues in our framework. We mainly follow cues in Liévin et al. [2022] as shown in Table 3. Since Liévin et al. [2022] focuses on medical domains, for general domains, we modify its cues by removing information specific to medical domains. We then generate self-explanations and perform ICL with them following our SELF-EXPLAIN framework. The final test results are shown in Table 4. We find no matter what cues are employed, ICL with self-generations elicited can all outperform using human-crafted CoTs, which demonstrates the robustness of SELF-EXPLAIN framework on the choice of cues.
Appendix B Effects of Generation Diversity
In this section, we examine the effects of generation diversity in SELF-EXPLAIN framework. By default, for each input of the exemplar, the LLM is prompted to generate five different ways of explanations to improve the generalization to test cases. At test time, one of five self-explanations is randomly chosen for an exemplar. To better understand the effects of this component in SELF-EXPLAIN, we therefore experiment with prompting different numbers of self-explanations for one exemplar input. Results are shown in Fig. 6. We find more diverse self-explanations to choose for an in-context exemplar can generally improve the test performance, while such improvement experiences diminishing return with further increased number of generations.
Appendix C Experiment Implementation Details for Sec. 6
We use MedCQA as testbed which offers human-crafted explanations for all the data. We use text-davinci-003 since it provides access to output probabilities. For the experiments on selection bias, we modify the correct options in in-context exemplars to make them the same in each case, e.g., all correct options are option A so that to give better control of influence of ICL. We append each option to the question to get the model’ confidence of the option token.
Appendix D Example self-explanation
D.1 MedQA
Question: A 22-year-old male marathon runner presents to the office with the complaint of right-sided rib pain when he runs long distances. Physical examination reveals normal heart and lung findings and an exhalation dysfunction at ribs 4-5 on the right. Which of the following muscles or muscle groups will be most useful in correcting this dysfunction utilizing a direct method? (A) anterior scalene (B) latissimus dorsi (C) pectoralis minor (D) quadratus lumborum
Self-explanation: The patient is a pregnant woman, and nitrofurantoin is considered safe for use during pregnancy. Other antibiotics, such as ciprofloxacin and doxycycline, are contraindicated in pregnancy due to potential harm to the fetus.
Human-crafted explanation: We refer to Wikipedia articles on medicine for help. Among the options, only pectoralis minor muscle origins from the outer surfaces of the 3rd to 5th ribs.
Question: A 44-year-old man comes to the office because of a 3-day history of sore throat, nonproductive cough, runny nose, and frontal headache. He says the headache is worse in the morning and ibuprofen does provide some relief. He has not had shortness of breath. Medical history is unremarkable. He takes no medications other than the ibuprofen for pain. Vital signs are temperature 37.4 C (99.4 F), pulse 88/min, respirations 18/min, and blood pressure 120/84 mm Hg. Examination of the nares shows erythematous mucous membranes. Examination of the throat shows erythema and follicular lymphoid hyperplasia on the posterior oropharynx. There is no palpable cervical adenopathy. Lungs are clear to auscultation. Which of the following is the most likely cause of this patient2̆019s symptoms? (A) Allergic rhinitis (B) Epstein-Barr virus (C) Mycoplasma pneumonia (D) Rhinovirus
Self-explanation: The presentation of a fussy infant who is regurgitating all of his feeds and has yellow vomit is consistent with pyloric stenosis, which is characterized by abnormal hypertrophy of the pylorus. This condition causes narrowing of the passage between the stomach and the small intestine, leading to difficulty in food passage and subsequent vomiting.
Human-crafted explanation: We refer to Wikipedia articles on medicine for help. The symptoms, especially the headache, suggest that the most likely cause is Rhinovirus. Epstein-Barr virus will cause swollen lymph nodes but there is no palpable cervical adenopathy. Lungs are clear to auscultation suggests it2̆019s not Mycoplasma pneumonia.
D.2 MedCQA
Question: Which of the following statements about Total Infra Venous Anaesthesia (TIVA) is true? (A) Causes More Renal Toxicity (B) Reduces Cerebral Metabolic Rate (C) Risk of Malignant Hyperthermia is high (D) Inhibits Hypoxic Pulmonary Vasconstriction
Self-explanation: By understanding the mechanism of action: Total Intravenous Anaesthesia (TIVA) involves the administration of intravenous drugs to induce and maintain anesthesia. One of the main drugs used in TIVA is propofol, which is known to reduce cerebral metabolic rate.
Human-crafted explanation: Total intravenous anaesthesia refers to a technique in general anaesthesia using a compination of agents given solely by IV route and in the absence of all inhalation agents. TIVA (Total intravenous anaesthesia) is produced by IV propoful. Propofol decreases the cerebral metabolic rate. Propofol does not impair renal function, does not trigger malignant hyperthermia and does not inhibit hypoxic pulmonary vasoconstriction.
Question: Shock causes: Choose the answer from A to D. A: Stagnant hypoxia. B: Anemic hypoxia. C: Hypoxic hypoxia. D: Histotoxic hypoxia.
Self-explanation: Shock is characterized by a decrease in blood pressure and impaired circulation. This can result in reduced blood flow to the tissues, leading to stagnant hypoxia.
Human-crafted explanation: Ans: A (Stagnant hypoxia)Ref: Pat GK, Hypoxia and Oxygen Therapy, In: Textbook of Medical Physiology, 2nd ed, Abuja Publishing House 2011: 94: 754-55
D.3 StrategyQA
Question: Would ramen be bad for someone with heart failure?
Self-explanation: Ramen can be bad for someone with heart failure due to its high sodium content. Sodium can cause fluid retention and increase blood pressure, which can put additional strain on the heart.
Human-crafted explanation: People with heart failure have to limit their sodium intake. Ramen is notorious for having incredibly high sodium levels.
Questions: Is Cape Town south of the Equator?
Self-explanation: By looking at a map: You can visually locate Cape Town on a map and observe that it is situated below the Equator, in the southern hemisphere.
Human-crafted explanation: Cape Town is an important city in South Africa. South Africa is located entirely south of the Equator.